OpenStudy (anonymous):

Msin(Bt)+Ncos(Bt)=squareroot(M^2+N^2)sin(Bt+C)

OpenStudy (anonymous):

Is this how it is supposed to look?

OpenStudy (anonymous):

\[M \sin Bt+N \cos Bt=\sqrt{M^2+N^2} \sin (Bt+C)\]

OpenStudy (anonymous):

Trying to solutions for the variable t, right?

OpenStudy (anonymous):

\[M \sin Bt+N \cos Bt=\sqrt{M^2+N^2}\div \sqrt{M^2+N^2}timesM \sin Bt+N \cos Bt\]

OpenStudy (anonymous):

I know I should also use the sum identity for the sine function. But, I'm stuck and have a blank on what to do next.

OpenStudy (anonymous):

t = -(1/B) cos^-1